# Deformable Beta Splatting

### v1

@@ 4.3. Beta Decomposition
By setting a mask based on the $b$ parameter, we can selectively isolate primitives based on their frequency contributions. We can also do a simple binary split based on a threshold (e.g. the mean of $b$ values across all primitives). Primitives with $b$ below the threshold primarily represent the scene’s solid geometry, with primitives of $b$ above the threashold provides high-frequency texture details.
@@
Since Beta Splatting integrates a learnable Phong Reflection Model (Phong, 1975), we can extracting diffuse component through taking only the base color $c_{0}$. Conversly, only enabling specular component gives the view-dependent specular lighting effects. For comparison, we only enable zero degree of SH for diffuse color while non zero degree for specular effects.
The results in Figure 6 illustrate the effectiveness of this approach. In the ”bicycle” scene, the black metal chair is completely separated from the white highlights, a result unattainable with Gaussian Splatting. Similarly, in the ”bonsai” scene, Beta Splatting successfully reconstructs the bright highlight component of the bonsai container, while Gaussian Splatting fails to capture such extreme view-dependent details.
@@ A.2. Inverse Splatting
This approach allows us to define a 2D kernel with the desired properties and ensure that a corresponding 3D kernel exists that satisfies the projection condition.
@@
Let’s consider the 2D kernel:
@@
with $\beta>0$ and $0\leq r\leq 1$. The derivative of $K(r)$ is:
@@
A.3. Requirements for the 2D Kernel Function
To ensure that the 2D kernel $K(r)$ corresponds to a valid 3D kernel $K(R)$ via the inverse Abel transform and satisfies the projection condition, the following requirements must be met:
@@ A.3. Requirements for the 2D Kernel Function
This approach allows us to design kernels with desired properties without needing to compute the 3D kernel explicitly.
Appendix B Kernel-Agnostic Distribution-Preserved Densification
Figure 7. Kernel-Agostic Densification illustration.
(32) $\alpha=o\cdot f(x)$
(33) $\alpha^{\prime}=1-(1-o^{\prime}\cdot f(x))^{N}$
(34) $o\cdot f(x)=1-(1-o^{\prime}\cdot f(x))^{N}$
Ignoring the kernel momentarily, the condition simplifies to:
(35) $o^{\prime}=1-\sqrt[N]{1-o},$
(36) $o^{\prime}=1-\sqrt[N]{1-o}\approx\frac{o}{N}$
Now consider the densified distribution:
we find:
This matches the original distribution:
(40) $\alpha=o\cdot f(x),$
up to first-order in $o$.
As illustrated in Figure 7, we can observe that the error decreases as opacity value becomes smaller. When opcacity gets close to 0.1, the two distributions are almost identical.
Appendix C Hyperparameters
Appendix D Additional Results
@@
This detailed per-scene analysis reveals that our method performs comparably to prior works on diffuse scenes, such as ”bicycle,” ”stump,” and ”flowers” from the Mip-NeRF 360 dataset. However, on scenes with complex specular reflections, such as ”counter,” Beta Splatting achieves significant improvements (more than 1 in PSNR) compared to 3DGS-MCMC, Zip-NeRF, and other methods. This highlights the robustness of our approach in capturing intricate light interactions in challenging scenes.

### v2

@@ 4.3. Beta Decomposition
By setting a mask based on the $b$ parameter, we can selectively isolate primitives based on their frequency contributions. We can also do a simple binary split based on a threshold (e.g. the mean of $b$ values across all primitives). Primitives with $b$ below the threshold primarily represent the scene’s solid geometry, with primitives of $b$ above the threshold provides high-frequency texture details.
@@
Since Beta Splatting integrates a learnable Phong Reflection Model (Phong, 1975), we can extract diffuse component through taking only the base color $c_{0}$. Conversely, only enabling specular component gives the view-dependent specular lighting effects. For comparison, we only enable zero degree of SH for diffuse color while non zero degree for specular effects.
The results in Figure 6 illustrate the effectiveness of this approach. In the ”bicycle” scene, the black bench is separated from the white highlights, a result unattainable with Gaussian Splatting. Similarly, in the ”bonsai” scene, Beta Splatting successfully reconstructs the bright highlight component of the bonsai container, while Gaussian Splatting fails to capture such extreme view-dependent details.
@@ A.2. Inverse Splatting
This guarantees that any suitably defined 2D kernel admits a multi-view-consistent 3D counterpart.
@@
Let’s take the Beta Kernel as an example:
@@
with $\beta>0$ and $0\leq r\leq 1$. The radial derivative of $K(r)$ is:
@@
This is the multi-view-consistent geometry we observe in 3D space by defining a 2D Beta Kernel in splatted billboards.
A.3. General Conditions for 2D Kernel Functions
To ensure that the 2D kernel $K(r)$ corresponds to a valid 3D kernel $K(R)$ via the inverse Abel transform and satisfies the projection condition, the following conditions must be met:
@@ A.3. General Conditions for 2D Kernel Functions
By enforcing the above conditions on any radial 2D function, the inverse Abel transform ensures the existence of a unique radial 3D kernel. Hence, one may prescribe any 2D splatting kernel and be guaranteed a corresponding, multi-view-consistent 3D geometry—without having to derive the 3D kernel in closed form.
@@
Appendix B Per-scene Results
Our per-scene evaluation on the Mip-NeRF 360 dataset shows that, on diffuse scenes (e.g., “bicycle,” “stump,” and “flowers”), Deformable Beta Splatting achieves the comparable performance of prior works, while on scenes with strong specular reflections (e.g., “counter”), it surpasses 3DGS-MCMC and Zip-NeRF by over 1 dB in PSNR. This gain demonstrates the ability of DBS to faithfully reconstruct intricate specular lighting in challenging scenarios.
Appendix C Kernel Comparison
Figure 7. Reconstruction quality on the Mip-NeRF 360 without view-dependent color-encodings. Deformable Beta Splatting (DBS) consistently outperforms 3D Gaussian Splatting (3DGS), confirming the stronger representation ability of the Beta kernel.
In addition to Table 3, we isolate the effect of the splatting kernel through an ablation study, disabling all view-dependent color terms (sh=0, sb=0) and comparing the standard Gaussian kernel (3DGS) to our deformable Beta kernel (DBS). Both models are trained on the Mip-NeRF 360 training set and evaluated on test views under MCMC densification schedules.
Figure 7 plots PSNR (left axis) and MSE (right axis) as a function of the total number of primitives. Although both kernels improve with more primitives, the Gaussian baseline plateaus around 1.5 M, while the Beta kernel continues to yield gains. Across the entire budget (0.1 M–5 M primitives), DBS consistently achieves higher PSNR and lower MSE, demonstrating that the deformable Beta kernel alone provides superior geometric modeling capacity.
Appendix D Kernel-Agnostic Distribution-Preserved Densification
D.1. Detailed Proof
Figure 8. Kernel-Agnostic Distribution-Preserved Densification illustration.
(31) $\alpha=o\cdot f(x)$
(32) $\alpha^{\prime}=1-(1-o^{\prime}\cdot f(x))^{N}$
(33) $o\cdot f(x)=1-(1-o^{\prime}\cdot f(x))^{N}$
Ignoring the kernel momentarily, the condition simplifies to:
(34) $o^{\prime}=1-\sqrt[N]{1-o},$
(35) $o^{\prime}=1-\sqrt[N]{1-o}\approx\frac{o}{N}$
Now consider the densified distribution:
we find:
This matches the original distribution:
(39) $\alpha=o\cdot f(x),$
up to first-order in $o$.
D.2. Validity of the Small Opacity Assumption
A natural question arises regarding the role of the small opacity prior used in our framework. At first glance, encouraging primitives to remain semi-transparent may appear counterintuitive: in surface-based representations, higher opacity is typically desirable to achieve clean geometry. However, our approach lies in Volumetric Density field. Volumetric methods—such as NeRF and 3DGS—fundamentally differ from surface‐based techniques. In a volumetric framework, each primitive contributes only a small opacity, and the final image emerges from the accumulation of many low‐opacity contributions. Within this paradigm, enforcing high per-primitive opacity can improve geometric fidelity and facilitate mesh extraction, but contradicts the core assumptions of smooth transmittance accumulation and degrades visual quality. Notable examples of this trade-off include NeuS (Wang et al., 2021) vs NeRF (Mildenhall et al., 2021) and SuGaR (Guédon and Lepetit, 2023) vs 3DGS (Kerbl et al., 2023). In contrast, applying a moderate opacity prior aligns with the volumetric principle, supports distribution-preserving densification, and further stabilizes the optimization process.
To quantify the impact of the opacity regularization, we analyze its effect both theoretically and empirically. In Figure 8, we visualize the original distribution (blue) and its densified distribution (red) as defined in Equation 33, across different kernel shapes. As opacity decreases, the discrepancy between the two distributions diminishes, with near-perfect alignment observed when the opacity approaches 0.1. This behavior supports our theoretical analysis: under the small-opacity regime, the densified distribution accurately preserves the original, regardless of the number of clones or the kernel function used.
$\lambda_{o}$ PSNR Opacity
0.000 27.81 0.60 $\pm$ 0.05
0.001 28.31 0.26 $\pm$ 0.05
0.005 28.61 0.14 $\pm$ 0.04
0.010 28.70 0.10 $\pm$ 0.03
0.050 28.69 0.04 $\pm$ 0.02
0.100 25.07 0.02 $\pm$ 0.02
Table 6. Mean PSNR and Opacity Distribution on the Mip-360 dataset for different $\lambda_{o}$ values.
Complementing this, we conduct an ablation study on the opacity regularization weight $\lambda_{o}$ across the Mip-NeRF 360 dataset. We report mean PSNR on held-out test views along with the mean and standard deviation of primitive opacities. As shown in Table 6, omitting the opacity term causes the densification step to naively clone primitives, resulting in over-saturated reconstructions and reducing visual quality. Introducing the opacity prior yields consistent gains in PSNR by reducing both the average opacity and its variance. However, excessively large $\lambda_{o}$ values overly suppress opacity and drive primitives toward excessive transparency, again causing damage to the quality of reconstruction. Empirically, we observe that setting $\lambda_{o}=0.01$ produces an average opacity around 0.1 and achieves peak PSNR, thereby validating our small opacity assumption.
[…]
3DGS 27.20 0.815 0.214 3.35 752.74 1
DBS‑full 28.75 0.845 0.179 3.11 356.04 2.11
DBS‑full (compressed) 28.60 0.840 0.182 3.11 63.48 41.70 11.86
DBS‑1M 28.42 0.831 0.205 1.00 114.44 6.58
DBS‑1M (compressed) 28.32 0.828 0.207 1.00 22.32 14.28 33.72
DBS‑490K 28.03 0.816 0.232 0.49 56.08 13.42
DBS‑490K (compressed) 27.80 0.806 0.237 0.49 11.37 9.15 66.20
Table 7. Quantitative evaluation of visual quality and model size on the Mip-360 dataset. More compression comparisons are available at 3DGS.zip (Bagdasarian et al., 2025).
Appendix E Compression
In Table 1 and Table 2, we configured DBS-full to use the similar number of primitives as vanilla 3DGS (3.35 M) for a fair comparison. Thanks to the MCMC framework, we can precisely control the target primitive count—so we also evaluate two ”pruned” variants, DBS-1M (1 M primitives) and DBS-490 (490 K primitives), to explore the trade-off between model size and visual quality. After training, we apply a post-training compression pipeline adapted from (Morgenstern et al., 2025) to further reduce each model’s on-disk footprint. Specifically, we firstly sort and reorder the parameters so that similar values are adjacent, maximizing the efficiency of lossless PNG compression. Next, we quantize most parameters from 32-bit floats down to 8-bit, allowing us to save them into PNG images. However, because we observed that reducing the precision of the 3D position to 8 bits causes a severe degradation in rendering quality, we preserve those values at full 32-bit precision, distributing them across four separate images, while all other parameters are stored as 8-bit PNGs.
We conduct these experiments on the Mip-360 dataset and report mean metrics across nine scenes in Table 7. As the model is pruned from 3.11 M primitives (DBS-full) to 1 M (DBS-1M) and then to 0.49 M (DBS-490K), we observe a graceful degradation in rendering fidelity: mean PSNR decreases from 28.75 dB to 28.42 dB to 28.03 dB, SSIM from 0.845 to 0.831 to 0.816, and LPIPS increases from 0.179 to 0.205 to 0.232. In parallel, the raw on-disk footprint contracts substantially—from 356 MB (2.11x relative to vanilla 3DGS) to 114 MB (6.58x) to 56 MB (13.42x). Applying uniform post-training PNG compression further amplifies these storage savings by approximately 5x–6x: final sizes become 63 MB, 22 MB, and 11 MB, corresponding to overall compression ratios of 11.9x, 33.7x, and 66.2x relative to 3DGS. Crucially, this additional compression incurs only minor quality losses (PSNR drops of 0.15 dB, 0.10 dB, and 0.23 dB) and remains computationally practical, requiring only 41.7 s, 14.3 s, and 9.2 s per scene, respectively. The naively compressed DBS-490K matches vanilla 3DGS’s visual quality in only $\sim$10 MB.
Appendix F Hyperparameters